OpenStudy (anonymous):

One x-intercept for a parabola is at the point (1,0). Use the factor method to find the other x-intercept for the parabola defined by this equation: y=4x2+4x−8 Separate the values with a comma.

OpenStudy (cwrw238):

one zeros of the equation is 1 which means that one factor is (x - 1)

OpenStudy (cwrw238):

so we have y = (x - a)(4x + a) = 4x^2 + 4x -8) can you find the values of a ?

OpenStudy (cwrw238):

* sorry its (x - 1)(4x + a) = 4x^2 + 4x - 8